• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Deur

    Log entry time 00:29:25 on October 6,2002

    Entry number 87406


    keyword=Shift summary

    16:51 Stop data taking for FPP straight through (run 1145). 470 k of events
    Dave meekins is here to investigate why our ion chambers trip when we have the optic carbon target in.
    16:58 5 uA beam for Dave's study. Target position is between two dummy targets (target encoder reads 2130747). Run 1146
    (target encoder reads 2151227). run 1147
    Encoder value for the the optic run position: 2489801
    17:18. Controled access to put in the FPP doors in (we will fix also q1 right arm that had tripped).
    Put the target on c12 optics
    17:55. Done with the access.
    We are going to take optics data.
    Put the collimator out for the right arm (left collimator already out)
    Ask for the beam back, as much current as possible, no raster
    Still no good beam available (tuning through compton chicane).
    The eP ion chamber is reading high dose (sometime it saturates at 10000 rads/h.).
    19:06 Q1 right tripped again. Asked for a controled access (we will wait for the tunning to be done).
    19:35 Finish tuning thru the compton chicane. We go in controled access.
    19:50 Q1 fixed. Back to beam permit.
    Still tuning.
    Problem with the bcm that is reading zero while beam was on.
    20:42 Beam is back but it still trips when I>1 uA. So the problem with the optic c12 target is not fixed. We move to BeO viewer to see if it is still a problem with the c12. Run with BeO is fine.
    -> called Dave Meekins
    We moved the target to a slightly different position: 2539801 but it does not help.
    We try 3000983 without success (dose on ion chambers is even worst).
    We took 2 short runs (1148 & 1149) at 1 uA (Huge rate: 75% of deadtime) to have a look at y_target to see if we understand the problem (Meekins suggested that something was hanging in front of it).
    We give up the optic data to take LH2 data (15 cm LH2 no radiator).
    21:53 LH2 ready: ask for 5 uA
    22:00 Beam cannot be delivered because the eP ion chamber trips. So mcc is going to tune the beam.
    22:15 The cooling flow from ESR is not very stable. As a consequence the density of the target is fluctuating a little bit. In addition mcc is not sure to be able to deliver us high intensity beam (due to eP ion chamber trips).
    So this is not good conditions for the linearity test so we drop it to move to physics data taking. We asked mcc to steer the beam better so they can deliver as high as they can.
    22:21 we have 5uA and mcc will try to give more. -> Start physics data taking
    We have 20 uA but the beam pos is bad: x=1.115-0.573 1.287+0.941
    we ask mcc to steer the beam better (they lower the beam down to 7 uA for the steering).
    We kept taking data (Ron Gilman said the elastic data was quiet insensitive to the beam position). The run where the beam position moved a lot (+/- 1mm in x and +/- 2mm in y) is run # 1150
    Beam position (absolute) is now: x=~0.2, y=~0.4
    23:03: beam trip. MCC is sending us tune beam now (we end the run : 1151 is a short run).
    23:45. Beam is back (first 5 uA and increasing to 18 uA. After that, beam steering becomes difficult so we will stay at 18-20 uA). Beam position is good (x=0.32, y=0.20) First run number is 1152. Prescalers have been modified for this run.